OpenStudy (anonymous):

Find all solutions in the interval [0, 2pi): Sin 3x = sin x

OpenStudy (anonymous):

sin(3x)=sin(2x+x)=sin(x)*cos(2x)+sin(2x)*cos(x)=sin(x)*cos(2x)+2sin(x)cos(x)*cos(x) sin(x)=0 is a solution so x=k*pi (I) wich k is an integer sor sinx not equal zero we have cos(2x)+2cos²(x)=1 since cos²(x)=(1+cos(2x))/2 we have : cos(2x)+1+cos(2x)=1->cos(2x)=0 -> x=k*pi/2+pi/4 (II)

OpenStudy (dinamix):

\[\sin^{-1} \sin(3x) = \sin^{-1} \sin(x)\] \[3x=x\] \[2x=0\] x=0\[x= \pi+k \pi \] @ganeshie8 could u check my answer i think i didnt made any mistake right ?